The data below contain the number of defects observed on each of 250 lcd screens. hw1_q4.csv

a) What proportion of the screens have more than 3 defects? Give your answer to three decimal places.

b)What propotion of the screens have at most 3 defects? Give your answer to three decimal places.

c)What proportion of the screens have fewer than 3 defects? Give your answer to three decimal places.

d)What proportion of the screens have exactly 3 defects? Give your answer to three decimal places.

e)What proportion of the screens have at least 3 defects? Give your answer to three decimal places.

defects
3
2
6
1
3
2
6
0
0
3
5
1
1
1
3
3
1
0
2
3
2
3
2
0
2
2
3
1
2
3
4
1
1
4
1
5
1
1
6
1
5
5
2
0
1
1
3
2
1
1
1
3
6
3
3
2
1
1
3
3
0
2
4
1
1
0
1
2
1
0
6
1
0
2
1
1
3
0
4
4
1
0
2
1
3
1
3
5
2
1
3
3
5
2
4
1
2
0
4
6
1
4
1
1
1
2
2
5
2
1
4
5
1
1
1
0
2
5
0
1
1
3
4
1
1
3
1
3
3
1
3
4
0
6
6
1
2
2
1
1
4
1
0
1
2
1
1
6
1
1
1
2
4
6
3
0
3
0
3
1
1
2
1
1
2
1
4
0
3
1
1
2
1
1
3
3
2
4
2
4
6
3
1
4
3
6
4
1
0
2
4
1
0
6
3
2
1
2
2
5
3
4
2
4
0
0
2
3
4
1
2
4
5
4
3
3
4
2
0
4
0
4
1
1
4
5
3
0
2
1
5
1
1
1
0
1
2
2
6
1
3
2
1
2
1
4
1
0
1
1
Answer #1

a) 56/250=0.224

b) 194/250=0.776

c) 154/250=0.616

d) 40/250=0.16

e) 96/250=0.384